Subject: Status of 3.5m, from Gillespie and Turner

From: Michael Strauss

Submitted: Mon, 25 Nov 2002 09:03:57 -0500 (EST)

Message number: 633 (previous: 632, next: 634 up: Index)

(Most of you have probably seen this by now, but here is the bad news,
if not.  Also see the <a href=>update by Russet McMillan</a>.


A sudden and unexpected mechanical problem in the 3.5-meter 
telescope's azimuth drives has occurred and will threaten the use of 
the telescope over the coming weeks.  It is entirely possible that 
the telescope will be shut down as early as next Monday, and not 
return to on-sky observing until about the 2nd week in December.  It 
is also quite possible that the telescope drive performance will
further degrade so seriously as to prevent operations at any time, thus
requiring an immediate shutdown.

The problem is in the drive boxes for the azimuth drive.  In azimuth, 
the telescope uses two DC motors whose drive capstans couple to the 
large drive surface, known as the drive tire, through a friction disk 
reduction box.  There are shafts and drive disks in these boxes that 
have required extensive maintenance and repairs over the lifetime of 
the telescope, usually on a time scale of every couple of years.  It 
appears now that something has caused the azimuth drive boxes to 
develop serious wear and slippage problems.  We usually schedule 
rebuilding of these boxes during summer shutdowns (e.g., the altitude 
drive box was rebuilt this past summer) before they fail, but the 
azimuth boxes have failed prematurely. We have spare parts sufficient 
to rebuild the damaged components, but the process involves a major 
shutdown of the telescope for a duration of roughly one to two weeks.

We were able to modify the system so that it could perform adequately 
last night, but the tracking performance of the telescope is marginal 
and probably not suitable for imaging programs.  The current symptom 
is a couple of arc-seconds wobble in azimuth during tracking, with a 
quasi-period of several seconds.  The frequency is fast enough that 
the Nasmyth and echelle guiders cannot not compensate adequately.  It 
is obvious that shafts and/or disks in the drive boxes are slipping 
or have been damaged--steps have been taken to enable continued 
limited use of the telescope by reducing slew speeds and disabling 
one of the azimuth motors.

We will attempt to operate the telescope in this degraded state over 
the weekend and then take a decision on Monday whether or not to shut 
down for repairs or continue to operate likewise until after the 
Thanksgiving holidays.  An updated message will be posted at that 
time, but users should be aware that their programs between now and 
mid-December are at some risk--it would be advisable for users with 
imaging programs to have a spectroscopic backup, and programs that 
minimize telescope slews are preferred.

An unrelated (one presumes!) and serious failure of the Peyton Hall 
computer systems in Princeton has probably disabled the apo35-general 
email list and will cause unfortunate delays in the distribution of 
this message.  Until it becomes available again, we are attempting to 
distribute this message via the institutional Users Cmte reps and 
Schedulers.  So, to those two groups, please forward this to your 
institutional user communities and otherwise distribute the
information as best you can.  Obviously, those with time scheduled 
3.5-m use coming up during the next few weeks have the greatest need 
to know, especially anyone planning to travel to APO to use the 

Ed Turner
Bruce Gillespie

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 633 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to